OpenStudy (liv33):

Graph the function: y = 9 - 2x

OpenStudy (karatechopper):

Well, can you identify the y-intercept and the slope in this equation?

OpenStudy (karatechopper):

Also, welcome to OpenStudy! :)

OpenStudy (karatechopper):

are you there @Liv33 lol

OpenStudy (justindrewbieber):

I Believe it is, (2,5)

OpenStudy (liv33):

I don't really know where to start with the problem lol. I came here cause I wanted to understand it :)

OpenStudy (karatechopper):

@Liv33 Do you know how to put that equation into slope-intercept form? y=mx+ b form?

OpenStudy (liv33):

vaguely

OpenStudy (karatechopper):

y = 9 - 2x In this case -2x is you mx and 9 is your b Now, can you place it in y=mx+b form for me?

OpenStudy (justindrewbieber):

yeah correct Karate an the answer is 2,5.

OpenStudy (liv33):

y = -2x +9

OpenStudy (karatechopper):

@JustinDrewBieber Liv is looking for a graph..I don't see where you are getting 2,5 from.

OpenStudy (karatechopper):

Good liv! Now, y=mx+b is the slope-intercept formula in which m=slope (rise/run) and b= the y-intercept So in this case, what is the y-intercept?

OpenStudy (liv33):

-2x + 9 ?

OpenStudy (karatechopper):

y=mx+b y=-2x+9 We are asking what the y-intercept (b) is, try again. :)

OpenStudy (liv33):

9 ?

OpenStudy (justindrewbieber):

uh, 18

OpenStudy (karatechopper):

Correct liv. @JustinDrewBieber Please do not interrupt the thread if you are not necessarily helping.

OpenStudy (karatechopper):

Right ok so if 9 is your y-intercept, how do you write that in coordinate form? (a,b)

OpenStudy (justindrewbieber):

Forgive My Mistakes Liv, I Was Confused. GoodLuck!

OpenStudy (liv33):

( ,9)

OpenStudy (karatechopper):

What would go in the space before the 9 lol

OpenStudy (liv33):

(-2, 9) ?

OpenStudy (liv33):

@karatechopper

OpenStudy (karatechopper):

Uhmm hm Ok so for an intercept, these points are graphed on their respected axes. So if asking for an x-intercept, you would say (x,0) if asked for a y-intercept, you would say (0,y). Make sense?

OpenStudy (liv33):

so (0, 9)

OpenStudy (karatechopper):

Correct. Now, can you graph that for me?

OpenStudy (liv33):



OpenStudy (karatechopper):

Good. Ok now, a slope is something that we refer to as rise over run. What is your slope (m) in y=-2x+9

OpenStudy (liv33):

-2

OpenStudy (karatechopper):

Looks good. Now if rise over run is your slope. then we hav to put this in fraction form, you agree that -2= -2/1 correct?

OpenStudy (liv33):

Yes

OpenStudy (karatechopper):

Alright, So if slope is rise/run and we have -2/1, then We are essentially rising by -2 (but in this case you will be going down by 2 because it is a negative value), and going to the right (if run is positive) 1. Do you understand that?

OpenStudy (liv33):

Yes, that makes sense

OpenStudy (sachintha):

I you plot it you will have a graph like this.